OpenStudy (anonymous):

Find the length of the radius of a sphere with a surface area of 314.16 ft².

OpenStudy (anonymous):

Recall that the area of a circle is: \[A=\pi r^2\]So, we'll plug in 314.16 ft^2 for the area and solve for r: \[314.16 ft^2 = \pi r^2\] \[100 ft^2 = r^2\] \[r=\sqrt{100 ft^2}=?\]

OpenStudy (anonymous):

r=10?

OpenStudy (anonymous):

Yep, 10 ft.

OpenStudy (bradely):

step by step answers posted here http://www.mathskey.com/question2answer/4145/find-the-length-of-the-radius-of-a-sphere-with-surface-area-of ask your pre-calculus homework questions at http://www.mathskey.com/question2answer/ … and get free math help. all the best